# Mutually-Aware Feature Learning for Few-Shot Object Counting

### v1

@@ 4.3 Comparison with the State-of-the-Art
To evaluate the model’s robustness against target confusion, assessments need to be conducted in multi-class scenes where diverse class objects coexist. Given that the FSC-147 dataset mainly consists of simple scenes with only target class objects, its limited multi-class scenes are not suitable for target confusion evaluation. Thus, we construct a multi-class subset of the FSC-147 (FSC-147-Multi), where images contain non-target objects more than $20~{}\%$ of the number of existing target class objects. Initially, we compare our model with state-of-the-art (SOTA) methods in a complex multi-class scenario, and then extend the evaluation to a simpler single-class scenario.
@@ 4.3.1 Evaluation on a multi-class scenario
We compare our method with SOTA methods on multi-class datasets: FSCD-LVIS and FSC-147-Multi, as shown in Table 2. For a fair comparison on FSCD-LVIS, we reproduce all baseline models using an image size of $512$ by utilizing official codes. Regarding Counting-DETR, we report its official performance. On the FSCD-LVIS, our method outperforms all baselines, showing an improvement of $11~{}\%$ in MAE and $4.6~{}\%$ in RMSE compared to the second-best performer. Similarly, on the FSC-147-Multi, our results demonstrate a significant performance gap compared to the SOTA method, showing a relative improvement of $26.30~{}\%$ in MAE and $24.63~{}\%$ in RMSE.
@@ 4.3.2 Evaluation on a single-class scenario
We further evaluate our method on the validation and test sets of the FSC-147 dataset, as shown in Table 3. Our method surpasses all baselines, showing an enhancement in MAEs of $12.9\%$ and $8.8\%$ for validation and test sets, respectively. It demonstrates the effectiveness of MAFEA even in a single-class scenario. In addition, we report the mean and standard deviation of results with $95\%$ confidence intervals over five runs. Notably, the test set’s RMSE exhibits a high standard deviation. To analyze a factor influencing test RMSE, we conduct evaluations by excluding extremely high-dense images with over 1000 instances. A total of 2 (over 1190) images are excluded from the test set. This exclusion leads to a notable improvement in LOCA’s RMSE, decreasing from $56.97$ to $32.85$ ($42.3\%$). Similarly, MAFEA’s performance dramatically improves with RMSE decreasing from $56.68$ to $34.35$ ($39.4\%$). This confirms the sensitivity of RMSE to errors in high-dense images and indicates the dependency of RMSE on how well the model fits extremely high-dense images.
4.3.3 Qualitative results
In Fig. 3, the 1st and 2nd rows illustrate qualitative results of the FSCD-LVIS. As shown in the 3rd to 6th columns, prior methods struggle to identify the target class in a query using exemplar images; they often count not only target objects but also non-target objects sharing similar scale or appearance properties. In contrast, our method excels in precisely distinguishing target objects based on the exemplar images, a success attributable to our mutually-aware feature learning. The 3rd and 4th rows show the results for the FSC-147. Our method makes more accurate predictions compared to other methods, especially on dense images.
@@ 4.3.3 Qualitative results
GMN Lu et al. (2019) $32.92$ $39.88$
FamNet Ranjan et al. (2021) $28.84$ $44.47$
RCAC Gong et al. (2022) $17.98$ $24.21$
BMNet+ Shi et al. (2022) $10.44$ $13.77$
SPDCN Lin et al. (2022) $18.15$ $21.61$
SAFECount You et al. (2023) $16.66$ $24.08$
LOCA Djukić et al. (2023) $9.97$ $12.51$
CSTrans Gao and Huang (2024) $20.84$ $24.64$
@@ 4.4 Cross-Dataset Generalization
We evaluate the generalization ability of our model on a car counting dataset, CARPK. To avoid overlap between the train and test sets, the tested models are pre-trained with FSC-147 by excluding its car category. The results are summarized in Table 4. Note that, we do not fine-tune our model on the CARPK. As reported, our method outperforms the current state-of-the-art methods. It demonstrates the robustness of our method in cross-dataset generalization.
@@ 4.5 Ablation Study
Table 6: Ablation study on Mutual Relation Modeling (MRM), Background Token (BT), and Target-Background Discriminitive (TBD) Loss in FSC-147-Multi. ‘ALL’ denotes the performance on the entire area of the image. ‘Target’ denotes performance within the area encompassing all target objects, and ‘Non-Target’ signifies performance in the complementary area.
@@ 4.5.1 Component-level analysis on multi-class scenario
Furthermore, we assess performance in the target and non-target regions to verify whether the models count only the target objects. This is imperative since the evaluation within the entire region may compensate for potential under-predictions in the target region by incorrect predictions in the non-target region. In the few-shot object counting (FSOC), each object is annotated only with its center point, which is not sufficient to estimate good boundaries between the target region and non-target region. To overcome this limitation, we expand each point annotation to encompass an area equivalent to the maximum size of exemplars. Consequently, the target region encompasses all target objects, while the non-target region covers the complementary area. We provide the visualization of the target region map and corresponding results in Figure 4. If there is no target confusion issue, the model’s prediction in the non-target region should be zero. Impressively, MRM, BT, and TBD bring substantial performance improvements in both target and non-target areas. The notable enhancement in the non-target area validates that the proposed components alleviate target confusion as intended.
4.5.2 Component-level analysis on single-class scenario
@@
$1$ $13.30$ $24.26$ $8.32$ $18.80$ $11.14$ $42.37$ $11.68$ $99.17$
$2$ $12.81$ $23.68$ $6.52$ $15.32$ $10.79$ $42.04$ $10.91$ $80.74$
$3$ $12.47$ $22.94$ $6.36$ $9.18$ $8.92$ $32.45$ $9.84$ $56.68$
@@ 4.5.3 The number of exemplars
Figure 5: Alignment Score (AS) map of exemplar features and background token.
@@
Figure 6: Additional Qualitative Results on FSC-147 dataset.
4.5.4 Influence of the background token
To validate whether the background token highlights the background region, we visualize the alignment score (AS) map in Fig. 5. We also compute the AS map for the exemplar features by summing up the scores of these features. As shown in the Fig. 5, the exemplar features activate only the target areas, while the background token highlights the background areas including the non-target class objects. These results show that our model can effectively differentiate between target objects and background regions.
4.6 Additional Qualitative Results
We show more qualitative results. As reported in Figure 6(a), we can observe that our model makes precise predictions from sparse to dense scenes for the novel classes. However, as seen in Figure 6(b), accurate predictions are challenging for images with over 1000 instances, which are very dense. This issue is not unique to MAFEA. Given the practical application of few-shot object counting models, we believe that having 1000 objects within a single image is an uncommon scenario.
@@ 5 Conclusion
In this work, we discover and solve the target confusion problem, which arises when multiple object classes coexist in the query image, resulting in inaccurate identification of the target objects. To settle this problem, we introduce a novel perspective, Mutually-Aware Feature Learning (MAFEA), which encourages interaction between query and exemplar features from the outset. By incorporating mutual relation modeling into the feature extractor, our model produces highly target-aware query features, facilitating the distinction between target and non-target objects. Additionally, the background token with Target-Background Discriminative (TBD) loss enables the model to effectively differentiate target objects from background features. We demonstrate the robustness of our method against the target confusion problem through evaluations in a complex multi-class scenario, such as FSCD-LVIS and FSC-147-Multi. Moreover, we validate its effectiveness even in a single-class scenario. Experimental results validate the unique merits of our proposed method by achieving state-of-the-art performances and demonstrating its effectiveness in addressing the target confusion problem.

### v2

@@ 4.3 Comparison with the State-of-the-Art
To evaluate the model’s robustness against target confusion, assessments need to be conducted in multi-class scenes where diverse class objects coexist. Given that the FSC-147 dataset mainly consists of simple scenes with only target class objects, its limited multi-class scenes are not suitable for target confusion evaluation. Thus, we construct a multi-class subset of the FSC-147 (FSC-147-Multi), where images contain non-target objects more than $20~\%$ of the number of existing target class objects. Initially, we compare our model with state-of-the-art (SOTA) methods in a complex multi-class scenario and then extend the evaluation to a simpler single-class scenario.
@@ 4.3.1 Evaluation on a multi-class scenario
We compare our method with SOTA methods on multi-class datasets: FSCD-LVIS and FSC-147-Multi, as shown in Table 2. On the FSCD-LVIS, our method outperforms all baselines, showing an improvement of $11~\%$ in MAE and $4.6~\%$ in RMSE compared to the second-best performer. On the FSC-147-Multi, our results demonstrate a significant performance gap compared to the SOTA methods in both zero- and few-shot settings. In the zero-shot setting, methods can be categorized into two types: exemplar-based (class name) and exemplar-free ($\times$). Exemplar-free methods, including ours, do not receive target class information and only count dominant class objects in the image. Despite this, they outperform exemplar-based methods, indicating that existing zero-shot methods face challenges with target confusion. In the few-shot setting, our method shows exceptional performance, thanks to its effective addressing of target confusion.
In Fig. 3, the 1st and 2nd rows illustrate qualitative results of the FSCD-LVIS. As shown in the 3rd to 6th columns, prior methods struggle to identify the target class in a query using exemplar images; they often miscount non-target objects that share similar scale or appearance with the target. In Fig. 4, we visualize results when different exemplars are provided in a multi-class scene. Despite the distinct appearances of the exemplar images, previous methods suffer from target confusion, localizing all object-like areas in the query image, regardless of the exemplar provided. In contrast, our method excels in precisely distinguishing target objects based on the exemplar images, a success attributable to our mutually-aware feature learning.
@@ 4.3.2 Evaluation on a single-class scenario
We further evaluate our method on validation and test sets of FSC-147, as shown in Table 3. Our method surpasses all baselines in both zero-shot and few-shot settings, demonstrating the effectiveness of MAFEA, even in a single-class scenario. As shown in Fig. 3, the 3rd and 4th rows show the results for the FSC-147. Our method makes more accurate predictions compared to other methods, especially on dense images.
@@
GMN GMN $32.92$ $39.88$
FamNet FamNet $28.84$ $44.47$
RCAC RCAC $17.98$ $24.21$
BMNet+ BMNet+ $10.44$ $13.77$
SPDCN SPDCN $18.15$ $21.61$
SAFECount SAFECount $16.66$ $24.08$
LOCA LOCA $9.97$ $12.51$
CSTrans cstrans $20.84$ $24.64$
@@
Figure 4: Qualitative results for a multi-class scenario. Each box in the query image is a box annotation for an exemplar, while the numbers in the pictures are the counting results. Best viewed with zoom-in.
@@ 4.4 Cross-Dataset Generalization
We evaluate the generalization ability of our model on a car counting dataset, CARPK. To avoid overlap between the train and test sets, the tested models are pre-trained with FSC-147 by excluding its car category. The results are summarized in Table 4. Note that we do not fine-tune our model on the CARPK. As reported, our method outperforms the current state-of-the-art methods. It demonstrates the robustness of our method in cross-dataset generalization.
@@ 4.5 Ablation Study
Table 6: Ablation study on Mutual Relation Modeling (MRM), Background Token (BT), and Target-Background Discriminitive (TBD) Loss in FSC-147-Multi.
@@ 4.5.1 Component-level analysis on multi-class scenario
Furthermore, we assess performance in the target and non-target regions to verify whether the models count only the target objects. This is imperative since the evaluation within the entire region may compensate for potential under-predictions in the target region by incorrect predictions in the non-target region. In the few-shot object counting (FSOC), each object is annotated only with its center point, which is insufficient to estimate good boundaries between the target and non-target region. To overcome this limitation, we expand each point annotation to encompass an area equivalent to the maximum size of exemplars. Consequently, the target region encompasses all target objects, while the non-target region covers the complementary area. We provide the visualization of the target region map and corresponding results in Figure 5. If there is no target confusion issue, the model’s prediction in the non-target region should be zero. Impressively, MRM, BT, and TBD bring substantial performance improvements in both target and non-target areas. The notable enhancement in the non-target area validates that the proposed components alleviate target confusion as intended.
@@
11 $13.30$ $24.26$ $8.32$ $18.80$ $11.14$ $42.37$ $11.68$ $99.17$
22 $12.81$ $23.68$ $6.52$ $15.32$ $10.79$ $42.04$ $10.91$ $80.74$
33 $12.47$ $22.94$ $6.36$ $9.18$ $8.92$ $32.45$ $9.84$ $56.68$
4.5.2 Component-level analysis on single-class scenario
@@ 4.5.3 The number of exemplars
Figure 6: Alignment Score (AS) map of exemplar features and background token. 2nd and 3rd columns are from model trained w/o TBD loss, and the 4th and 5th columns are from model trained w/TBD loss.
4.5.4 Influence of the background token
To confirm that BT highlights the background region, including non-target objects, we visualize the alignment score (AS) map in Fig. 6. AS map is obtained by summing up the alignment scores of BT or exemplar features. BT alone fails to activate the non-target objects because there is no mechanism enforcing its interaction with the query’s background. Our TBD loss directly aligns exemplars and BT with the query’s target and non-target features. As depicted in Fig. 6, BT operates as intended only when used with TBD loss.
Additionally, we conduct an ablation study on the number of BTs in Fig. 7. While removing BT leads to degraded performance, the quantity of tokens has a minor impact on the results. BT aims to prevent undesired interactions between the query’s non-target features and the exemplar features. Therefore, the presence of the BT, rather than its quantity, truly matters.
4.6 Complexity Comparison
We compare the computational complexity and the number of parameters in Table 9. Although our model uses more parameters than CNN-based models, its computational cost is similar or lower. Compared to ViT-based models like CounTR, our model uses fewer parameters.
4.7 Additional Qualitative Results
We present more qualitative results. As shown in Figure 8(a), our model accurately predicts counts across sparse and dense scenes for novel classes. However, our method sometimes confuses the target objects when non-target objects have a similar appearance or size to the target class, as seen in the 1st row of Figure 8(b). Moreover, our method may mistakenly count both classes when the exemplar image contains multiple classes, as shown in the 2nd row. Additionally, in the 3rd to 5th rows, our model struggles with images that contain incredibly high object densities, especially those with more than 1000 instances.
Figure 7: Results w.r.t the number of background tokens.
Table 9: Computational complexity and the number of parameters.
Method Encoder Param (M) GFLOPS (G)
SAFECount CNN $32$ $366$
LOCA CNN $35$ $80$
CounTR ViT $100$ $91$
MAFEA (Ours) ViT $93$ $122$
@@ 4.7 Additional Qualitative Results
Figure 8: Success cases and failure cases on FSC-147 dataset.
@@ 5 Conclusion
In this work, we highlight and address the target confusion problem in multi-class scenarios. We conjecture that target confusion arises because existing methods encode query and exemplar features independently, without any explicit interaction during feature extraction. This lack of interaction causes insufficient target awareness, making it difficult for the model to distinguish between target and non-target objects in multi-class scenes. To settle this problem, we designed the model to distinguish between target and non-target objects by producing target-aware features through early interaction in the feature extraction process. The comprehensive results validate that our proposed method effectively mitigates target confusion. However, as discussed in Sec. 4.7, despite its robustness, our method often fails to identify the correct target class when objects from visually similar classes co-exist. To overcome this limitation, future work could explore fine-grained classification techniques, which are specifically designed to capture subtle differences between visually similar classes. Incorporating such techniques could improve the model’s ability to distinguish visually similar classes, thereby enhancing its real-world performance.
6 Appendix
Figure 9: Details about the Relation Learner. Input images $I^{Q}$ and $I^{E}$ are encoded into features $z^{Q}$ and $z^{E}$. The Object Prototype Extraction (OPE) module combines shape properties from n bounding boxes with appearance properties from exemplar features $z^{E}$ to generate n object prototypes. $z^{Q}$ is depth-wise correlated (*) with n object prototypes, resulting in n similarity maps. The response map is then obtained by computing the per-element maximum of n similarity maps and upsampled by the decoder to the final density map.
6.1 Additional Details
As described in Sec. 3, our research focus lies in the feature extractor with Mutual Relation Modeling (MRM). We adopt the same relation learner and decoder as LOCA, so we do not claim contribution in this regard. In this section, we provide details about the relation learner, as illustrated in Figure 9. The relation learner, inspired by LOCA, consists of the Object Prototype Extraction (OPE) and Prototype Matching (PM) modules. The sole difference lies in the OPE modules, where LOCA obtains appearance features from ROI pooling in the query feature ($z^{Q})$ using box annotations, while we leverage exemplar features ($z^{E}$) from the MRM encoder. The spatial size (s) of exemplar features is set to $3$, consistent with LOCA.